OpenStudy (jmartinez638):

Geometric sequence with 12 as its second term, and 324 as its fifth. What is the third term?

OpenStudy (anonymous):

\[a,ar,ar^2,ar^3,ar^4\]

OpenStudy (anonymous):

\[ar=12\] and \[ar^4=324\] divide to solve for \(r\)

OpenStudy (anonymous):

if that is not clear, divide means \[\frac{ar^4}{ar}=r^3=\frac{324}{12}=27\] telling you that \[r^3=27\]

OpenStudy (jmartinez638):

That makes sense. Thanks!

OpenStudy (anonymous):

yw hope you got that \(r=3\) and the third term is \[12\times 3\]